# PIR Over Wireless Channels: Achieving Privacy With Public Responses

### v1

@@ Abstract
This paper addresses the problem of private information retrieval (PIR) over an additive white Gaussian noise (AWGN) channel, considering the channel is public. In such settings, each server can eavesdrop on the channel, potentially compromising the user’s privacy. Previous works suggested joint coding–PIR schemes, ignoring the fact that communication over a practical wireless channel is public. To address this gap, we present a novel joint wiretap–PIR coding scheme that leverages lattice codes to exploit the channel’s additive properties. This scheme integrates wiretap coding and private retrieval techniques into a unified framework.
@@ I Introduction
The PIR problem has recently attracted the attention of information theory society, motivated by the challenge of characterizing the fundamental limits of this problem. The classical PIR capacity, with noiseless and orthogonal channels, is given by ${C_{PIR}=(1-1/N)(1-(1/N)^{M})}$ in [1]. This key finding encourages the information theory community to explore further variations of this issue. Since then, many extensions and generalizations of the problem have been examined. For instance, robust PIR considers scenarios where some databases may fail to respond, and $T$-private PIR ensures privacy even when any $T$ out of the $N$ databases collude, as studied in [2]. Another variation is PIR from byzantine databases, which accounts for the possibility of up to $B$ databases providing erroneous responses, either intentionally or unintentionally [3].
@@
While the PIR problem has been extensively studied over the past decades, most research has focused on simple communication channels characterized by orthogonal and noiseless links between the user and the database. The case of PIR over noisy orthogonal channels (NPIR) was examined in [4], which demonstrated that the channel coding required to mitigate channel errors is ”almost separable” from the retrieval scheme, depending only on the agreed traffic ratio. Additionally, [4] explored PIR over various types of MAC and showed that, unlike NPIR, the channel coding and retrieval scheme is not optimal in general. Building on this foundation, the authors in [5, 6, 7] investigated PIR over a block-fading Gaussian MAC.
@@
In those papers, the authors did not refer to the fact that the channel model, which represents a practical wireless channel, is publicly available, hence making it possible for the servers to overhear the channel, which may cause a privacy leakage. This paper tackles the secrecy gap by utilizing methods developed for wiretap coding. The main idea is to prevent the servers from retaining any information related to the user from the observed signal.
@@
In [8], the author extends the classical PIR problem to address scenarios involving colluding servers and eavesdroppers. This extended problem introduces additional privacy and secrecy constraints to the conventional PIR setting. Specifically, the user aims to retrieve a message without revealing the desired message index to any set of $T$ colluding, and an eavesdropper who can listen to the queries and answers of any $E$ servers but is prevented from learning any information about the messages. This problem is related to ours in the sense that it adds the secrecy constraint, and it can provide some insights. However, unlike [8], which assumes noiseless and orthogonal channels, our model assumes a MAC channel between the servers and the user, which presents a significant difference. The server could not collude in our model but can serve as an eavesdropper.
This paper is a direct continuation of [5], which tackles the problem of PIR over block–fading AWGN channel by considering that servers can also listen to the same channel used for their transmissions. We provide a joint wiretap-privacy-channel coding scheme to handle this. The resulting achievable rate is compared to the full channel capacity without any constraint when sending a single message. This comparison shows the redundancy of the suggested joint scheme, one has to tolerate to promise privacy.
@@ II-A Notational Conventions
Throughout the paper, we use boldface lowercase to refer to vectors, e.g., $\mathbf{h}\in\mathbb{R}^{L}$, and boldface uppercase to refer to matrices, e.g., $\mathbf{H}\in\mathbb{R}^{M\times L}$. For a vector $\mathbf{h}$, we write ${\|\mathbf{h}\|}$ for its Euclidean norm, i.e. ${\|\mathbf{h}\|}\triangleq\sqrt{\sum_{i}h_{i}^{2}}$. We denote by $\mathbf{e}_{i}$ the unit vector with $1$ at the $i$th entry and zero elsewhere. All logarithms in this paper are natural logarithms, and rates are in nats.
@@ II-B System Model and Problem Statement
We assume the servers receive the queries through independent control channels and cannot access each other’s queries or answers, a common assumption in the literature.
The $k$th server responds to its query with a message (or codeword) $\mathbf{x}_{k}(i)$ of size $n$, where $n$ is fixed for all responses. The reason for this is that we follow the usual Gaussian MAC setup in the literature [9], where a codeword is transmitted during $n$ channel uses. The $k$th server generates an answer based on the received query $Q_{k}(\theta)$ and the stored messages $W_{1}^{M}$. The answer, denoted by $\mathbf{x}_{k}(i)$, is a deterministic function of the messages and the query. We have
@@
Because the server’s responses are transmitted over a public channel, each server can monitor other servers’ communications, potentially compromising user privacy. To ensure privacy, the server should not be able to infer any information about the desired message index given the queries, the entire database, and the received signal $\mathbf{w}$ at the curious server antenna. Thus, for each server $j$ the following privacy constraint must be satisfied,
$\begin{split}\text{[Secure - Privacy]}&\\ &\lim_{n\rightarrow\infty}I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w})=0,\hskip 5.69046pt\forall j\in\{1,...,N\}.\end{split}$ (4)
The databases are linked to the user via a block-memoryless fading AWGN channel (Figure 1). In this setup, the channel remains constant throughout the transmission of codewords of size $n$, and each block is independent of the others. Thus, over a transmission of $n$ symbols, the user and the curious server observe a noisy linear combination of the transmitted signals,
$\mathbf{y}=\sum_{k=1}^{N}h_{k}\mathbf{x}_{k}+\mathbf{n}_{y},\quad\mathbf{w}=\sum_{k=1}^{N}g_{k}\mathbf{x}_{k}+\mathbf{n}_{w}$ (5)
Here, $h_{k}\thicksim\mathcal{N}(0,1)$ and $g_{k}\thicksim\mathcal{N}(0,1)$ represent the real channel coefficients of the servers-user and servers-server channel, respectively, and $\mathbf{n}_{y}\thicksim\mathcal{N}(0,\sigma_{y}^{2}\mathbf{I}^{n\times n})$,$\mathbf{n}_{w}\thicksim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are i.i.d. Gaussian noise observed by the user and the curious server, respectively.
@@
We assume there is channel state information at the transmitter (CSIT) together with statistical information of the curious server, i.e., $\{h_{k}\}_{k=1}^{N}$,$\{g_{k}\}_{k=1}^{N}$ and $\sigma_{w}^{2}$ are known at the transmitter. Upon receiving the mixed response $\mathbf{y}$ from all the databases, the user decodes the required message $W_{i}$. Let $\widehat{W_{i}}$ denote the decoded message at the user and define the error probability for decoding a message as follows,
@@
Remark 1:
Note that the [Secure - Privacy] requirement, as defined above, is a special case that arises from the AWGN channel characteristic and is different from the usual conventions in the PIR filled. In our case, we will see that ensuring the Secure-Privacy constraint requires a strong secrecy notion from the wiretap channel.
Figure 1: System model of two servers connected to a user through an AWGN MAC channel while server $2$ is also aware of the transmission over the channel.
@@ II-C Performance Metric
The PIR rate over MAC is defined as the total number of desired bits divided by the total number of channel uses [5, 7]. Specifically,
@@
Neglecting the privacy constraint, our model matches with Multiple Input Single Output (MISO) channel with per-antenna power constraint model [10] hence this model will serve as an upper bound to our PIR capacity rate. In such a model, each transmit antenna has a separate power budget yet can fully cooperate with other antennas. The MISO sum capacity with per-antenna power constraint and fixed channel coefficients, which are globally known, is given by
$C_{SR}^{MISO}=\frac{1}{2}\log\left(1+P\left(\sum_{k=1}^{N}|h_{k}|\right)^{2}\right).$ (7)
We note that the MAC sum capacity is lower than $C_{SR}^{MISO}$ since, in the latter, the transmit antenna can fully cooperate with each other.
@@ II-D Lattice Code and Lattice Gaussian Distribution
for all $\mathbf{x}\in\mathbb{R}^{n}$. Note that $f_{\sigma,\Lambda}$ restricted to the quotient $\mathbb{R}^{n}/\Lambda$ 111the set of all closets which in our case equivalent to the Voronoi region is a probability density.
@@ III Main results
The following theorem presents an achievable strong secrecy PIR rate for the AWGN channel without fading, i.e., all the fading coefficients equal $1$. In our suggested scheme, we use the results provided in [14], which propose a wiretap lattice code that achieves strong secrecy and semantic security over the Gaussian wiretap channel (GWC). Their scheme achieves the secrecy capacity of the GWC up to a constant gap of $\frac{1}{2}$ nat. The first result is for $N=2$ servers.
@@
For the $N=2$ databases public AWGN channel, the following strong secrecy PIR rate is achievable,
$R<\frac{1}{2}\log^{+}\left(\min\left\{\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{w}},SNR_{y}\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{y}}\right\}\right)-\frac{1}{2}$ (14)
where $SNR_{y}\frac{P}{\sigma^{2}_{y}}$, and $SNR_{w}=\frac{P}{\sigma^{2}_{w}}$.
@@
Consider the PIR problem with $N\geq 2$ databases over a public AWGN. Then, for any non-empty subsets of databases $\mathcal{S}_{1},\mathcal{S}_{2}$ satisfying ${\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}$, $\mathcal{S}_{1}\cup\mathcal{S}_{2}\subseteq\{1,...,N\}$, the following PIR rate is achievable,
$R<\frac{1}{2}\log^{+}\left(\min\left\{\frac{\frac{1}{2}+SNR_{y}^{\prime}}{1+SNR_{w}^{\prime}},SNR_{y}^{\prime}\frac{\frac{1}{2}+SNR_{y}^{\prime}}{1+SNR_{y}^{\prime}}\right\}\right)-\frac{1}{2}$ (15)
where $SNR_{y}^{\prime}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{y}}$, $SNR_{z}^{\prime}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{w}}$.
Accordingly, the achievable rate scaling laws with the number of servers $N$ and the power $P$ are the same as those for the wiretap channel. The next theorem considers the PIR over public AWGN block–fading channel,
@@
$R<\frac{1}{2}\log^{+}\left(\min_{i}\left\{\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{w}_{i}}},SNR_{\tilde{y}}\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{y}}}\right\}\right)-\frac{1}{2}$ (16)
@@ IV A Joint PIR Scheme Over Public AWGN MISO channel
This section presents a PIR over the public AWGN MISO scheme.
@@ IV-A A Joint Scheme for Secure PIR Over Non-Fading AWGN MAC With $N=2$ Servers
We start by analyzing the simple case with $N=2$ servers, as depicted in Figure 1. We assume that reciprocity between the channels. Therefore, it suffices to assume that only one server is eavesdropping on the channel. Let $\mathbf{y}$ and $\mathbf{z}$ denote the received inputs at the user and the malicious server (server number $2$), respectively. Over a transmission of $n$ symbols, the received input can be described as,
@@
$\mathbf{w}(i)=\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{n}_{w}$
where $\mathbf{n}_{y}\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbf{I}^{n\times n})$ and $\mathbf{n}_{w}\sim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are the noises at the legitimate receiver and the eavesdropper, respectively. Assume, without loss of generality, that server $2$ can cancel its transmission. Thus, server $2$ obtains the following
$\mathbf{z}_{1}(i)=\mathbf{x}_{1}+\mathbf{n}_{w}.$
Note that the index $i$ in the received input denotes the private index of the desired message.
In the following, we provide an achievability scheme that proves Theorem 1.

### v2

@@ Abstract
In this paper, we address the problem of Private Information Retrieval (PIR) over a public Additive White Gaussian Noise (AWGN) channel. In such a setup, the server’s responses are visible to other servers. Thus, a curious server can listen to the other responses, compromising the user’s privacy. Indeed, previous works on PIR over a shared medium assumed the servers cannot instantaneously listen to other responses. To address this gap, we present a novel randomized lattice – PIR coding scheme that jointly codes for privacy, channel noise, and curious servers which may listen to other responses. We demonstrate that a positive PIR rate is achievable even in cases where the channel to the curious server is stronger than the channel to the user.
@@ I Introduction
The PIR problem has recently attracted the attention of information theory society, motivated by the challenge of characterizing the fundamental limits of this problem. The classical PIR capacity, with noiseless and orthogonal channels, is given by ${C_{PIR}=(1-1/N)(1-(1/N)^{M})}$ in [1]. This key finding encourages the information theory community to explore further variations of this issue, tailored to the specific challenges arising naturally in real life. Many extensions and generalizations of the problem have been examined. For instance, robust PIR considers scenarios where some databases may fail to respond, and $T$-private PIR ensures privacy even when any $T$ out of the $N$ databases collude, as studied in [2]. Another variation is PIR from byzantine databases, which accounts for the possibility of up to $B$ databases providing erroneous responses, either intentionally or unintentionally [3].
@@
While the PIR problem has been extensively studied over the past decade, most research has focused on simple communication channels characterized by orthogonal and noiseless links between the user and the database. The case of PIR over noisy orthogonal channels (NPIR) was examined in [4], which demonstrated that the channel coding required to mitigate channel errors is ”almost separable” from the retrieval scheme, depending only on the agreed traffic ratio. Additionally, [4] explored PIR over various types of MAC and showed that, unlike NPIR, the channel coding and retrieval scheme is not optimal in general. Building on this foundation, the authors in [5, 6, 7] investigated PIR over a block-fading Gaussian MAC.
@@
In these papers, the authors did not refer to the fact that the channel model, which represents a practical wireless channel, is publicly available, hence making it possible for a curious server to overhear other responses, which may cause a privacy leakage. This paper aims to maintain privacy under the assumption that any of the servers can listen to the channel. The main goal is to prevent the servers from gaining any information about the identity of the user’s desired message from their observed signal.
@@
In [8], the author extends the classical PIR problem to address scenarios involving colluding servers and eavesdroppers. This extended problem introduces additional privacy and secrecy constraints to the conventional PIR setting. Specifically, the user aims to retrieve a message without revealing the desired message index to any set of $T$ colluding, and an eavesdropper who can listen to the queries and answers of any $E$ servers but is prevented from learning any information about the messages. This related problem indeed adds a secrecy constraint, yet it assumes noiseless and orthogonal channels. Our model assumes a Gaussian MAC channel between the servers and the user, which presents a significant difference.
To address these challenges, we propose a joint PIR-channel coding scheme that exploits the additive nature of the AWGN MAC while engineering the responses such that no server can infer the identity of the desired message. Simultaneously, the responses, which ”sum up over the air,” enable the user to decode the desired message. Additionally, the integration of wiretap coset coding ensures that any curious server is unable to decode the identity from the overheard responses, thereby preserving privacy.
Section II outlines the system model and provides the necessary preliminaries. Section III presents the achievable PIR rates for different variations of the proposed model. Section IV details the PIR achievable schemes and provides proofs for the Theorems introduced in Section III.
@@ II-A Notational Conventions
Throughout the paper, we use boldface lowercase to refer to vectors, e.g., $\mathbf{h}\in\mathbb{R}^{L}$. For a vector $\mathbf{h}$, we write ${\|\mathbf{h}\|}$ for its Euclidean norm, i.e. ${\|\mathbf{h}\|}\triangleq\sqrt{\sum_{i}h_{i}^{2}}$. All logarithms in this paper are natural logarithms, and rates are in nats.
@@ II-B System Model and Problem Statement
We assume the servers receive the queries through independent control channels and cannot access each other’s queries or answers [^1] This assumption is a common assumption in the literature and is justified by the fact that queries are low-rate messages, allowing encryption to be implemented without incurring significant overhead.
The $k$th server generates an answer based on the received query $Q_{k}(\theta)$ and the stored messages $W_{1}^{M}$. The answer, denoted by $\mathbf{x}_{k}(i)$, is of size $n$, where $n$ is fixed for all responses,222 This is because we follow the usual Gaussian MAC setup in the literature [9], where a codeword is transmitted during $n$ channel uses.. The answer is a deterministic function of the messages and the query. We have
@@
Because the server’s responses are transmitted over a public channel, a byzantine server could monitor the communication within the channel, potentially compromising user privacy (Figure 1). To ensure privacy, each server should not be able to infer any information about the desired message index given the queries, the entire database, and the received signal $\mathbf{w}_{j}$ at the $j$-th server antenna. Thus, for each server $j$ the following privacy constraint must be satisfied,
$\begin{split}\text{[Privacy - under public responses]}&\\ &\lim_{n\rightarrow\infty}I(\theta;Q_{j}(\theta),W_{1}^{M},\mathbf{w}_{j})=0,\hskip 5.69046pt\forall j\in\{1,...,N\}.\end{split}$ (4)
Note that the privacy requirement (4), as defined above, is a special case arising from all servers sharing the same communication channel. The main difference lies in the curious server’s knowledge of $\mathbf{w}_{j}$, which represents the received signal at its antenna. Additionally, it is important to highlight that our privacy requirement is asymptotic, as it relies on the strong secrecy notion to ensure no information leakage from $\mathbf{w}_{j}$.
The databases are linked to the user via a block-memoryless fading AWGN channel (Figure 1). In this setup, the channel remains constant throughout the transmission of codewords of size $n$, and each block is independent of the others. Thus, over a transmission of $n$ symbols, the user and the $j$-th server observe a noisy linear combination of the transmitted signals,
$\mathbf{y}=\sum_{k=1}^{N}h_{k}\mathbf{x}_{k}+\mathbf{n}_{y},\quad\mathbf{w}_{j}=\sum_{k=1}^{N}g_{j,k}\mathbf{x}_{k}+\mathbf{n}_{j,w}$ (5)
Here, $h_{k}\thicksim\mathcal{N}(0,1)$ and $g_{j,k}\thicksim\mathcal{N}(0,1)$ represent the real channel coefficients of the servers-user and servers-server channel, respectively, and $\mathbf{n}_{y}\thicksim\mathcal{N}(0,\sigma_{y}^{2}\mathbf{I}^{n\times n})$, $\mathbf{n}_{j,w}\thicksim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are i.i.d. Gaussian noise observed by the user and the curious server, respectively.
@@
We assume there is full channel state information at the transmitter (CSIT) and together with statistical information of the curious server, i.e., $\{h_{k}\}_{k=1}^{N}$,$\{g_{j,k}\}_{k=1}^{N}$ for any $j$, and $\sigma_{w}^{2}$ are known. We also assume reciprocity between the channels, meaning that the forward and reverse channels between the servers and the user exhibit the same properties and the same channel coefficients. Under the CSIT assumption, the curious server can either cancel its own transmitted message or remain idle while passively listening. Consequently, if the index of the malicious server is $j$, it obtains
$\mathbf{z_{j}}=\mathbf{w}_{j}-g_{j,j}\mathbf{x}_{j}=\sum_{k=1,k\neq j}^{N}g_{j,k}\mathbf{x}_{k}+\mathbf{n}_{j,w}.$ (6)
Upon receiving the mixed response $\mathbf{y}$ from all the databases, the user decodes the required message $W_{i}$. Let $\widehat{W_{i}}$ denote the decoded message at the user and define the error probability for decoding a message as follows,
@@
Figure 1: System model with two servers connected to a user via an AWGN MAC channel. Server $1$ acts as a malicious entity, monitoring the communication and having access to $w_{1}$.
@@ II-C Performance Metric
For a joint PIR-coding scheme over MAC, we define the PIR rate as the total number of desired bits divided by the total number of channel uses [5, 7]. Specifically,
@@
Neglecting the privacy constraint, our model matches with Gaussian MAC with per-antenna power constraint model [9], as the servers cannot explicitly cooperate but may implicitly do so through the queries together with the assumption of full CSIT. An upper bound for this model can be established by having all antennas transmit the same symbol simultaneously, effectively reducing the system to an AWGN channel with a single antenna and a total power equal to the sum of the individual powers. Consequently, this setup serves as an upper bound for the PIR capacity rate in our model.
$C_{AWGN}^{MAC}=\frac{1}{2}\log\left(1+\frac{P_{total}}{\sigma^{2}}\right).$ (8)
@@ II-D Lattice Code and Lattice Gaussian Distribution
for all $\mathbf{x}\in\mathbb{R}^{n}$. Note that $f_{\sigma,\Lambda}$ restricted to the quotient $\mathbb{R}^{n}/\Lambda$ 333the set of all cosets which in our case equivalent to the Voronoi region is a probability density.
@@ III Main results
The following theorem presents an achievable PIR rate for the AWGN channel with public responses and without fading, i.e., all the fading coefficients equal $1$. The first result is for $N=2$ servers.
@@
For the $N=2$ databases public AWGN channel, the following PIR rate is achievable,
$R=\frac{1}{2}\log\left(\min\left\{\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{w}},SNR_{y}\frac{\frac{1}{2}+SNR_{y}}{1+SNR_{y}}\right\}\right)-\frac{1}{2}$ (15)
where $SNR_{y}=\frac{P}{\sigma^{2}_{y}}$, and $SNR_{w}=\frac{P}{\sigma^{2}_{w}}$.
@@
Consider the PIR problem with $N\geq 2$ databases over a public AWGN channel. Then, for any non-empty subsets of databases $\mathcal{S}_{1},\mathcal{S}_{2}$ satisfying ${\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}$, $\mathcal{S}_{1}\cup\mathcal{S}_{2}\subseteq\{1,...,N\}$, the following PIR rate is achievable,
$R=\frac{1}{2}\log\left(\min\left\{\frac{\frac{1}{2}+SNR_{y}^{\prime}}{1+SNR_{w}^{\prime}},SNR_{y}^{\prime}\frac{\frac{1}{2}+SNR_{y}^{\prime}}{1+SNR_{y}^{\prime}}\right\}\right)-\frac{1}{2}$ (16)
where $SNR_{y}^{\prime}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{y}}$, $SNR_{w}^{\prime}=\left\lfloor{\frac{N}{2}}\right\rfloor^{2}\frac{P}{\sigma^{2}_{w}}$.
Accordingly, the achievable rate scaling laws with the number of servers $N$ and the power $P$ are the same as those for the AWGN wiretap channel. The next theorem considers the PIR over public AWGN block–fading channel,
@@
$R=\frac{1}{2}\log\left(\min_{i}\left\{\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{w}_{i}}},SNR_{\tilde{y}}\frac{\frac{1}{2}+SNR_{\tilde{y}}}{1+SNR_{\tilde{y}}}\right\}\right)-\frac{1}{2}$ (17)
@@ IV A Joint PIR Scheme Over Public AWGN MISO channel
This section presents the schemes and the proofs for a PIR over the public AWGN MISO scheme.
@@ IV-A A Joint Scheme for Secure PIR Over Non-Fading AWGN MAC With $N=2$ Servers
Consider the simple case with $N=2$ servers, as depicted in Figure 1. To ensure privacy constraints, focusing on the server with the best channel conditions is sufficient. In this two-server scenario, channel reciprocity is assumed; thus, without loss of generality, we designate server $2$ as the malicious server. Let $\mathbf{y}$ and $\mathbf{z}_{2}$ represent the received inputs at the user and the malicious server, respectively. Over a transmission of $n$ symbols, the received input can be described as,
@@
$\mathbf{z}_{2}(i)=\mathbf{x}_{1}+\mathbf{n}_{j,w}$
where $\mathbf{n}_{y}\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbf{I}^{n\times n})$ and $\mathbf{n}_{w}\sim\mathcal{N}(0,\sigma_{w}^{2}\mathbf{I}^{n\times n})$ are the noises at the legitimate receiver and the curious server, respectively (6),(5). Note that the index $i$ in the received input denotes the private index of the desired message.
In the following, we provide an achievability scheme that proves Theorem 1. In our suggested scheme, we use the results provided in [13], which propose a wiretap lattice code that achieves strong secrecy and semantic security over the Gaussian wiretap channel (GWC). Their scheme achieves the secrecy capacity of the GWC up to a constant gap of $\frac{1}{2}$ nat.